OpenStudy (anonymous):

find the prinipal root of sqrt{a^4b^2}

OpenStudy (anonymous):

\[\sqrt{a^4b^2}\]

OpenStudy (stamp):

\[\sqrt{a^4b^2}=(a^4b^2)^{1/2}\]

OpenStudy (anonymous):

oh boy ,, that doesn't even make sense to me

OpenStudy (stamp):

square root is the same as taking the 1/2 power

OpenStudy (mertsj):